OpenStudy (anonymous):

if it takes 2/5 of an hour to do 2/3 of the work, how long does it take to do all the work?

OpenStudy (ineedhelplz):

Welcome to openstudy! :)

OpenStudy (isaidavila):

Establish a ratio 2/5/x=2/3/1

OpenStudy (welshfella):

2/3 work takes 2/5 hours so all ... .. .. 2/5 / 2/3 = 2/5 * 3/2 hours